# Audio-Language Models for Audio-Centric Tasks: A survey

### v1

@@ V-B1 Multi-task Tuning Models
Instruction tuning is another one of the current mainstream methods, which aims to provide multiple audio tasks with unified instance construction formats using natural language. This allows a single model to handle multiple tasks without explicit fine-tuning on specific tasks but by enhancing its instruction-following capability [133]. SpeechGPT [8] incorporates discrete speech representations into a large language model framework, enabling intrinsic cross-modal conversational abilities. SpeechGPT is built upon SpeechInstruct, a pioneering large-scale cross-modal speech instruction dataset, and employs a sophisticated three-stage training strategy: modality-adaptation pre-training, cross-modal instruction fine-tuning, and chain-of-modality instruction fine-tuning. This approach allows SpeechGPT to perceive and generate multi-modal content effectively, demonstrating impressive capabilities in following cross-modal human instructions and showcasing the potential for handling multiple modalities within a single model. LLark [129] and MusiLingo [128] are instruction-tuned multimodal model designed specifically for music tasks, from classification and regression to more complex captioning and reasoning challenges and has demonstrated impressive capabilities in zero-shot tasks. SALMONN [119] conduct a three-stage cross-modal training method of pre-training, instruction tuning and an additional activation tuning stage proposed to resolve the issue of over-fitting to the speech recognition and audio captioning tasks in instruction tuning. SALMONN demonstrates impressive capabilities across zero-shot and downstream tasks in environmental sound, speech, and music, and introduce two novel tasks: audio-based storytelling and speech audio co-reasoning. LTU [24] aims to leverage the comprehension abilities of LLM to enable an audio foundation model to listen, think, and understand. To train LTU, they created a new OpenAQA-5M dataset consisting of closed-ended and open-ended diverse (audio, question, answer) tuples, using a perception-to-understanding curriculum with an autoregressive training framework. LTU demonstrates strong performance and generalization ability on conventional audio tasks such as classification and captioning, as well as emerging audio reasoning and compression abilities that are absent in existing audio models. GAMA [121] integrates multiple types of audio representations using a custom Audio Q-Former and a multi-layer aggregator, capturing diverse aspects from semantic generalization to surface-level audio properties. They introduced CompA-R and CompA-R-test, a human-labeled dataset for open-ended audio question answering requiring complex reasoning. Fine-tuned on CompA-R, GAMA demonstrates strong performance in complex reasoning and instruction-following tasks.
@@
In-context learning enables LLMs to quickly adapt to specific tasks using a few examples, allowing them to predict unseen input labels without additional parameter updates. In the transfer of ALMs, the goal is to learn and enhance vocabulary from the context of prompt texts without performing backpropagation [122], thereby unlocking the capabilities of frozen LLMs for a wide range of audio understanding and generation tasks [134]. Voicebox [125] is a non-autoregressive flow-matching model trained to infill speech based on audio context and text, utilizing over 50,000 hours of unfiltered or enhanced speech data for training. Voicebox can perform various tasks through in-context learning, such as mono or cross-lingual zero-shot TTS, noise removal, content editing, style conversion, and diverse sample generation, achieving high fidelity and efficiency in speech synthesis and editing. SALM [122] equipped with multitask and in-context learning capabilities. SALM achieving comparable performance to task-specific models through instruction tuning for ASR and AST. Additionally, a speech-supervised in-context training method is proposed to bridge the gap between LLM training and downstream speech tasks, further enhancing the in-context learning ability of speech-to-text models. Audio Flamingo [120] supports instruction tuning on interleaved audio-text pairs, thereby excelling in robust multi-turn dialogue capabilities. By incorporating in-context learning, it achieves quick adaptation to unseen tasks. The model attains state-of-the-art results on several close-ended and open-ended audio understanding tasks without the need for task-specific fine-tuning.
@@ V-B2 Audio-Language Agent Systems
SpeechAgents [131] is a multi-modal, multi-agent system designed to simulate human communication. The agents collaborate in a distributed and parallel manner to formulate decisions and execute corresponding actions, significantly enhancing work efficiency and effectiveness. A Multi-Agent Tuning technique was proposed, allowing the system to handle complex, multi-party communications more effectively. Notably, the system shows excellent scalability, performing well even with up to 25 agents. This capability makes SpeechAgents particularly suitable for tasks such as drama creation and audio novel generation. MusicAgent [132] is a system designed to integrate and organize various music processing tasks. The method employed by MusicAgent consists of a comprehensive toolset for collecting music-related tools from diverse sources, and an autonomous workflow powered by LLMs that organizes these tools, decomposes user requests into multiple sub-tasks automatically, and invokes the appropriate music tools to fulfill these requirements. MusicAgent is capable of handling a wide range of music processing tasks and aims to bridge the gap between tools from different sources by unifying data formats (text, MIDI, ABC notation, audio), thereby enabling seamless collaboration among tools on different platforms. AudioGPT [130] is an innovative multi-modal system that integrates ChatGPT with audio foundation models and input/output interfaces to process complex audio information and conduct spoken conversations. Its method involves four key stages: modality transformation, task analysis, model assignment, and response generation. AudioGPT demonstrates proficiency in various audio-related tasks, including speech processing, music analysis and generation, and sound effect understanding and creation.
SpeechAgents [131] is a multimodal and multi-agent system that proposes a multi-agent tuning technique. The agents collaborate in a distributed and parallel manner to make decisions and execute actions, thereby significantly improving work efficiency and effectiveness. With excellent scalability, SpeechAgents is particularly suitable for tasks such as drama creation and audio novel generation. MusicAgent [132] is a system designed to integrate and organize various music processing tasks. Driven by LLMs, it automatically decomposes user requests into multiple sub-tasks and invokes the appropriate music tools to fulfill these requirements. MusicAgent can handle a wide range of music processing tasks and aims to bridge the gap between tools from different sources by unifying data formats (text, MIDI, ABC notation, audio), enabling seamless collaboration among tools on different platforms. AudioGPT [130] is an innovative multimodal system that integrates ChatGPT with audio foundation models and input/output interfaces to process complex audio information and conduct spoken conversations. AudioGPT demonstrates proficiency in various audio-related tasks, including speech processing, music analysis and generation, and sound effect understanding and creation.
@@ VI Datasets and Benchmarks
Figure 11: Illustration of two types of audio-language datasets: audio-text paired datasets and audio question answering datasets.
@@ VI-A Audio-Text Paired Datasets
Audio-text paired datasets consist of audio recordings paired with corresponding text, such as captions (description of audios), transcriptions (for speech datasets) or translations (transcriptions in another language of what the speaker says). For natural language can describe multiple events in the audio more freely and in greater detail, audio-caption datasets providing more information for representation pre-training. Datasets composed of transcriptions and translations are more often used for training models on specific tasks. Commonly used audio-text paired datasets for training ALMs are listed in Table IV.
@@
AudioCaps [19] Audio AudioSet 57K Human
Sound-VECaps [135] AudioSet 1.6M Generated
AudioGrounding [136] AudioSet 5K Generated
AudioSetCaps [137] AudioSet, Youtube-8M, VGGSound 6M Generated
LAION-Audio-630K [22] AudioSet, FreeSound, BBC Sound Effects, etc. 630K -
WavCaps [40] AudioSet, FreeSound, BBC Sound Effects, SoundBible 400K Generated
Auto-ACD [138] AudioSet, VGGSound 1.5M Generated
Clotho [20] FreeSound 30K Human
AudioDiffCaps [139] FreeSound 29K Generated
augmented FSD50K [140] FreeSound 51K Online
WavText5k [141] BigSoundBank, SoundBible 5K Online
SoundDescs [15] BBC Sound Fffects 33K Online
MACS [142] TAU Urban Acoustic Scenes 17K Human
AudioCaption [10] Internal Collected 29K Human
EMOSpeech [143] Speech Internal Collected 31K Generated+Human
SpeechCraf [144] AISHELL-3, Zhvoice, GigasSpeech-m, LibriTTs-R 2.1M Generated
MusicCaps [90] Music AudioSet 5.5K Human
AnimalSpeak [145] Bioacoustic iNaturalist, Xeno-Canto, Animal Sound Archive, etc. 1.1M Generated
@@ VI-B Audio Question Answering Datasets
Audio Question Answering Datasets are designed for training and evaluating models that can answer questions about audio content. These datasets are structured as triplets of audio, question, and answer. Initially, such datasets were specifically created for Audio Question Answering (AQA) tasks, constructing diverse questions to train models on how to answer similar queries. In recent years, AQA tasks have been increasingly utilized as a proxy task for model learning perception and reasoning about audio [146]. They can serve as a format that integrates multiple audio tasks [24] and can also be used to guide models in better following human instructions [8]. We have listed several popular open-source QA datasets in Table V.
@@
Datasets Audio Source Pairs Text Source
@@
DAQA [146] AudioSet 600K
AudioSetCaps-QA [137] AudioSet, Youtube-8M, VGGSound 18.4M
SpeechInstruct [8] Gigaspeech, Common Voice, LibriSpeech 9M
ClothoAQA [21] FreeSound 36K
mClothoAQA [147] 287K
OpenAQA-5M [24] AudioSet, VGGSound, Freesound, Sound Bible, etc. 5.6M
CompA-R [26] AudioSet 200K
Audio Dialogues [148] AudioSet 163.8K
MusicQA [149] AudioSet,TagATune 70K
@@ VI-C Benchmarks
Benchmarks can be categorized into three types: task-specific, cross-task, and audio instruction benchmarks. Task-specific benchmarks like SoundDescs [15], CompA [26], and ADU-Bench [150] evaluate models on particular tasks such as ATR and audio dialogue. Cross-task benchmarks such as ARCH [151] and MMAU [152] test generalization across multiple domains and tasks, emphasizing diverse training data. Other speech-related cross-task benchmarks include SUPERB [153], Dynamic-SUPERB [154], LeBenchmark [155], LAPE [156], VoiceBench [157], as well as music domain benchmarks like MusicBench [91] and MuChoMusic [158]. Audio instruction benchmarks like AIR-Bench [159] and AudioBench [160] assess the abilities of LALMs in understanding audio signals and following instructions, using a unified human-computer conversation format to cover various speech and audio tasks.

### v2

@@ V-B1 Multi-task Tuning Models
Instruction tuning is another one of the current mainstream methods, which aims to provide multiple audio tasks with unified instance construction formats using natural language. This allows a single model to handle multiple tasks without explicit fine-tuning on specific tasks but by enhancing its instruction-following capability [liu2024instructiontuning]. SpeechGPT [zhang2023speechgpt] incorporates discrete speech representations into a large language model framework, enabling intrinsic cross-modal conversational abilities. SpeechGPT is built upon SpeechInstruct, a pioneering large-scale cross-modal speech instruction dataset, and employs a sophisticated three-stage training strategy: modality-adaptation pre-training, cross-modal instruction fine-tuning, and chain-of-modality instruction fine-tuning. This approach allows SpeechGPT to perceive and generate multi-modal content effectively, demonstrating impressive capabilities in following cross-modal human instructions and showcasing the potential for handling multiple modalities within a single model. LLark [gardner2023llark] and MusiLingo [deng2024musilingo] are instruction-tuned multimodal model designed specifically for music tasks, from classification and regression to more complex captioning and reasoning challenges and has demonstrated impressive capabilities in zero-shot tasks. SALMONN [tangsalmonn] employs a three-stage training process: pre-training, instruction tuning, and an additional activation tuning stage designed to mitigate overfitting to speech recognition and audio captioning tasks. It demonstrates strong zero-shot and downstream performance across environmental sound, speech, and music, while introducing two novel tasks: audio-based storytelling and speech-audio co-reasoning. LTU [gong2024ltu] integrates LLM comprehension with audio foundation models for listening, thinking, and understanding. Trained on the OpenAQA-5M dataset consisting of closed-ended and open-ended diverse (audio, question, answer) tuples, using a perception-to-understanding curriculum. It excels in conventional tasks such as classification and captioning, as well as emerging audio reasoning and compression abilities that are absent in existing audio models. GAMA [ghosh2024gama] integrates multiple types of audio representations using a custom Audio Q-Former and a multi-layer aggregator, capturing diverse aspects from semantic generalization to surface-level audio properties. They introduced CompA-R and CompA-R-test, a human-labeled dataset for open-ended audio question answering requiring complex reasoning. Fine-tuned on CompA-R, GAMA demonstrates strong performance in complex reasoning and instruction-following tasks.
@@
In-context learning enables LLMs to quickly adapt to specific tasks using a few examples, allowing them to predict unseen input labels without additional parameter updates. In the transfer of ALMs, the goal is to learn and enhance vocabulary from the context of prompt texts without performing backpropagation [Chen2024SALM], thereby unlocking the capabilities of frozen LLMs for a wide range of audio understanding and generation tasks [yang2024uniaudio1.5]. Voicebox [NIPS2023Voicebox] is a non-autoregressive flow-matching model trained to infill speech based on audio context and text, utilizing over 50,000 hours of unfiltered or enhanced speech data for training. Voicebox can perform various tasks through in-context learning, such as mono or cross-lingual zero-shot TTS, noise removal, content editing, style conversion, and diverse sample generation, achieving high fidelity and efficiency in speech synthesis and editing. SALM [Chen2024SALM] equipped with multitask and in-context learning capabilities. SALM achieving comparable performance to task-specific models through instruction tuning for ASR and AST. Additionally, a speech-supervised in-context training method is proposed to bridge the gap between LLM training and downstream speech tasks, further enhancing the in-context learning ability of speech-to-text models. Audio Flamingo [kong2024audioflamingo] achieves leading performance on various audio understanding tasks without requiring task-specific fine-tuning, supporting instruction tuning for robust multi-turn dialogue and demonstrating rapid adaptation through in-context learning. Building on this foundation, Audio Flamingo 2 extends these capabilities to long-form audio understanding and introduces the LongAudioBench benchmark [ghosh2025audioflamingo2]. The recently developed Audio Flamingo 3 [goel2025flamingo3] further expands both training data and learning curricula, achieving breakthrough advancements in reasoning and understanding across speech, sound, and music domains.
@@ V-B2 Audio-Language Agent Systems
SpeechAgents [zhang2024speechagents] is a multimodal and multi-agent system that proposes a multi-agent tuning technique. The agents collaborate in a distributed and parallel manner to make decisions and execute actions, thereby significantly improving work efficiency and effectiveness. With excellent scalability, SpeechAgents is particularly suitable for tasks such as drama creation and audio novel generation. MusicAgent [yu2023musicagent] is a system designed to integrate and organize various music processing tasks. Driven by LLMs, it automatically decomposes user requests into multiple sub-tasks and invokes the appropriate music tools to fulfill these requirements. MusicAgent can handle a wide range of music processing tasks and aims to bridge the gap between tools from different sources by unifying data formats (text, MIDI, ABC notation, audio), enabling seamless collaboration among tools on different platforms. AudioGPT [huang2024audiogpt] is an innovative multimodal system that integrates ChatGPT with audio foundation models and input/output interfaces to process complex audio information and conduct spoken conversations. AudioGPT demonstrates proficiency in various audio-related tasks, including speech processing, music analysis and generation, and sound effect understanding and creation.
Notably, while Tables II and III document architectural relationships between pre-trained models and downstream tasks, systematic evaluations comparing pre-training methods for specific applications remain limited. This gap stems from rapidly evolving task paradigms and ALMs’ multi-component complexity, which complicates generalizable model selection [elizalde2024CLAP2]. (For representative evaluations on commonly assessed tasks like AC and ATR, see Section VII-B.) We therefore recommend experimental evaluation within specific application contexts to identify suitable pre-trained models.
Beyond supervised fine-tuning (SFT) paradigms discussed above, reinforcement learning (RL) and preference optimization are emerging as new directions for ALM downstream transfer. RL enables learning from evaluative feedback via cumulative reward maximization, showing potential to surpass SFT in audio reasoning [li2025RL, wu2025audiothinker]. Direct preference optimization (DPO) offers a lightweight alternative, learning from comparative judgments without explicit reward modeling [huang2025step], and proves effective for subjective qualities like naturalness and musicality [broukhim2025prefersurvey].
@@ VI Datasets and Benchmarks
Figure 11: Illustration of two types of audio-language datasets.
@@ VI-A Audio-Text Paired Datasets
Audio-text paired datasets contain audio recordings paired with corresponding text, including captions (describing audio content), transcriptions (for speech), or translations (transcriptions in other languages). Natural language descriptions capture multiple audio events more freely and in detail, making audio-caption datasets particularly valuable for representation pre-training. Transcriptions and translations are typically used for task-specific model training. Table IV summarizes commonly used audio-text paired datasets for ALM training.
@@
AudioCaps [kim2019audiocaps] Audio AudioSet 57K Human
Sound-VECaps [yuan2025soundvecaps] AudioSet 1.6M Generated
AudioGrounding [xu2021AudioGrounding] AudioSet 5K Generated
AudioSetCaps [bai2025audiosetcaps] AudioSet, Youtube-8M, VGGSound 6M Generated
LAION-Audio-630K [wu2023large] AudioSet, FreeSound, BBC Sound Effects, etc. 630K -
WavCaps [mei2024wavcaps] AudioSet, FreeSound, BBC Sound Effects, SoundBible 400K Generated
Auto-ACD [sun2024auto] AudioSet, VGGSound 1.5M Generated
Clotho [drossos2020clotho] FreeSound 30K Human
AudioDiffCaps [Takeuchi2023AudioDiffCaps] FreeSound 29K Generated
augmented FSD50K [Liu2024decasesep] FreeSound 51K Online
WavText5k [deshmukh2022wavtext] BigSoundBank, SoundBible 5K Online
SoundDescs [koepke2022benchmarks] BBC Sound Fffects 33K Online
MACS [Martin2021bMACS] TAU Urban Acoustic Scenes 17K Human
AudioCaption [wu2019audio] Internal Collected 29K Human
EMOSpeech [bian2024emospeech] Speech Internal Collected 31K Generated+Human
SpeechCraf [jin2024speechcraft] AISHELL-3, Zhvoice, GigasSpeech-m, LibriTTs-R 2.1M Generated
MusicCaps [agostinelli2023musiclm] Music AudioSet 5.5K Human
AnimalSpeak [robinson2024bioacoustics] Bioacoustic iNaturalist, Xeno-Canto, Animal Sound Archive, etc. 1.1M Generated
[…]
@@ VI-B Audio Question Answering Datasets
Audio Question Answering Datasets are designed for training and evaluating models that can answer questions about audio content. These datasets are structured as triplets of audio, question, and answer. Initially, such datasets were specifically created for Audio Question Answering (AQA) tasks, constructing diverse questions to train models on how to answer similar queries. In recent years, AQA tasks have been increasingly utilized as a proxy task for model learning perception and reasoning about audio [fayek2020temporal]. They can serve as a format that integrates multiple audio tasks [gong2024ltu, Kuan2025ICASSP] and can also be used to guide models in better following human instructions [zhang2023speechgpt]. We have listed several popular open-source QA datasets in Table V. Furthermore, data annotation is shifting from manual labeling toward synthetic and weakly-supervised approaches, including LLM-generated audio captions and self-supervised pseudo-labels (see “Text source” columns in Tables IV and V, where “generated” denotes model-produced text). For detailed dataset analyses, we refer readers to dedicated audio-language dataset surveys [wijngaard2024ald_survey].
@@
[…]
@@
DAQA [fayek2020temporal] AudioSet 600K
AudioSetCaps-QA [bai2025audiosetcaps] AudioSet, Youtube-8M, VGGSound 18.4M
SpeechInstruct [zhang2023speechgpt] Gigaspeech, Common Voice, LibriSpeech 9M
ClothoAQA [lipping2022clotho] FreeSound 36K
mClothoAQA [behera23_interspeech] 287K
OpenAQA-5M [gong2024ltu] AudioSet, VGGSound, Freesound, Sound Bible, etc. 5.6M
CompA-R [ghosh2024compa] AudioSet 200K
Audio Dialogues [goel2024Dialogues] AudioSet 163.8K
MusicQA [liu2024music] AudioSet,TagATune 70K
[…]
@@ VI-C Benchmarks
Benchmarks for ALMs fall into three categories: task-specific, cross-task, and audio instruction. Task-specific benchmarks such as SoundDescs [koepke2022benchmarks], CompA [ghosh2024compa], and ADU-Bench [gao2024ADUbench] evaluate performance on defined tasks including audio-text retrieval and audio dialogue. Cross-task benchmarks like ARCH [la2024ARCH], MMAU [sakshi2025mmau], SSEU-bench [yin2025SSEUbench] and AQUA-bench [kuan2026aqua] assess model generalization across domains, requiring diverse training data. This category also includes speech-related cross-task benchmarks [yang2021superb, Huang2024Dynamic-Superb, evain2021LeBenchmark, Ghosh2022LAPE, chen2024voicebench] and music-specific ones [melechovsky2024mustango, weck2024muchomusic]. Audio instruction benchmarks including AIR-Bench [yang2024air] and AudioBench [wang2025audiobench] test LALMs’ audio understanding and instruction-following abilities through unified conversational formats. Recent benchmark studies have begun exploring additional dimensions to measure ALMs’ capabilities, with Audio Jailbreak [song2025ajailbreak] and LongAudioBench [ghosh2025audioflamingo2] focusing specifically on adversarial robustness and long-audio processing capabilities, respectively.